OpenStudy (anonymous):

find the midpoint of the following mid- point segment with endpoints (3,2) and (-5, -6) ? what is the formula also?

OpenStudy (anonymous):

Find the slope first. plug in 3,2 in the equation y-y1=m(x-x1)

OpenStudy (anonymous):

that gives you the formula!

OpenStudy (anonymous):

thx!!

OpenStudy (anonymous):

Midpoint means half means average of coordinate: x = (3 -5) /2 = ....? y = (2 -6) /2 = ....? Can you continue, also don't forget to learn the Midpoint formula!

OpenStudy (anonymous):

yea i just went and learned it ;D thank you fro yur help!!

OpenStudy (anonymous):

I'm happy as long as somebody learn something :)

OpenStudy (anonymous):

hahahaha yud make a good teacher!!

OpenStudy (anonymous):

I'm really busy today and somehow feel so good =)

OpenStudy (anonymous):

;D